Find the particular solution of the equation



e^2tdy/dx=4



Given that y=5, t=0

1
Expert's answer
2022-03-31T02:42:07-0400

"e^{2t}dy\/dt=4"

"dy\/4=dt\/e^{2t}"

"\frac{y}{4}=\frac{-1}{2e^{2t}}+C"

y=5, t=0

C=5/4+1/2=7/4

"y=\frac{-2}{e^{2t}}+7"